hooperl:

A city determines that a planned community must have at least 4 acres of developed and open space, and the difference between the number of developed acres, y, and the number of open acres, x, can be no more than 1. Which graph represents the system of inequalities for this scenario? x + y ≥ 4 y – x ≤ 1 On a coordinate plane, 2 solid straight lines are shown. The first line has a positive slope and goes through (negative 1, 0) and (0, 1). Everything above and to the left of the line is shaded. The second line has a negative slope and goes through (0, 4) and (4, 0). Everything to the left of the line is shaded. On a coordinate plane, 2 solid straight lines are shown. The first line has a positive slope and goes through (negative 1, 0) and (0, 1). Everything below and to the right of the line is shaded. The second line has a negative slope and goes through (0, 4) and (4, 0). Everything above and to the right of the line is shaded. On a coordinate plane, 2 solid straight lines are shown. The first line has a positive slope and goes through (negative 4, 0) and (0, 4). Everything to the left of the line is shaded. The second line has a negative slope and goes through (0, 1) and (1, 0). Everything to the left of the line is shaded. On a coordinate plane, 2 solid straight lines are shown. The first line has a positive slope and goes through (negative 4, 0) and (0, 4). Everything to the right of the line is shaded. The second line has a negative slope and goes through (0, 1) and (1, 0). Everything to the right of the line is shaded.

jhonyy9:

x + y ≥ 4 y – x ≤ 1 like a first step x +y = 4 -x +y = 1 0 2y = 5 y = 5/2 x +5/2 = 4 2x +5 = 8 2x = 3 x = 3/2  any idea now ? @xxDeppressionxx

jhonyy9:

@Laylalyssa

jhonyy9:

using this table you can draw the graph more easy of these inequality